# Large deviations in non-Markovian stochastic epidemics

### v1

@@ Abstract
We develop a framework for non-Markovian SIR and SIS models beyond mean field, utilizing the continuous-time random walk formalism. Using a gamma distribution for the infection and recovery inter-event times as a test case, we derive asymptotical late-time master equations with effective memory kernels and obtain analytical predictions for the final outbreak size distribution in the SIR model, and quasistationary distribution and disease lifetime in the SIS model. We show that increasing memory can greatly widen the outbreak size distribution and reduce the disease lifetime. We also show that rescaled Markovian models fail to capture fluctuations in the non-Markovian case. Overall, our findings, confirmed against numerical simulations, demonstrate that memory strongly shapes epidemic dynamics and paves the way for extending such analyses to structured populations.
@@
Introduction. The challenge of modeling and predicting pandemics has motivated extensive development of deterministic and stochastic epidemiological models [1, 2, 3, 4, 5, 6, 7, 8]. Their importance was underscored during COVID-19 [9, 10], when forecasts guided policies worldwide. These models typically divide populations into disease-status compartments. The SIS model [11] with susceptible and infected compartments, describes infections with short-lived immunity (e.g., influenza, or the common cold), while the SIR model [8] adds a recovered compartment to capture long-term immunity, which is relevant for diseases such as measles, smallpox, polio and COVID-19. Generalized models include reinfection (SIRS), an exposed compartment (SEIR) [6], time-varying rates [12, 13], or structured populations, where network-based approaches [14, 15, 16, 17, 18, 19] capture heterogeneity in population structure. Mean-field quantities have been computed for both models: the mean outbreak size and epidemic threshold in the SIR model, and endemic state and relaxation dynamics in the SIS model [11, 5, 3, 14]. Stochastic dynamics have also been explored, including the outbreak-size distribution for the SIR model and quasistationary distribution (QSD) and disease mean time to extinction (MTE) for the SIS model [20, 21, 13, 22, 23, 22, 24, 25, 26, 27, 28, 29, 30].
A key limitation of most approaches is the Markovian assumption [7, 31, 32]. In reality, infectious periods are often non-exponential, better described by log-normal, gamma, or Weibull distributions [33, 34, 35]. In structured populations the interactions may be diffusion-limited rather than reaction-limited, and cause heterogeneous, non-Markovian, infection times. Thus, infection and recovery are more realistically non-Markovian, but analytical methods developed for Markovian systems are generally ill-suited to handle such processes.
@@
In the realm of epidemics, many studies have addressed non-Markovian processes [36, 37, 38, 39, 40, 41, 42, 43, 44, 32, 45, 46, 47, 48], mostly focusing on mean-field dynamics [39] and network generalizations [38, 42, 43, 44, 32, 45, 46]. In addition, Starnini et al. [48] showed that such dynamics can be reproduced by rescaling Markovian rates [49]. On the simulation side, Boguñá et al. [41] proposed a generalized Gillespie algorithm, later refined into the more efficient and accurate Laplace Gillespie method by Masuda et al. [47]. Stochastic non-Markovian epidemics have also been studied in both SIR [36, 37, 40, 46] and SIS [38] models, but a systematic analysis of how non-Markovianity and demographic noise jointly shape large deviations is still lacking.
Recent progress has extended non-Markovian frameworks to chemical and biological systems. Aquino & Dentz [50] developed a formalism based on continuous-time random walk (CTRW) with non-exponential waiting times (WTs), later adapted in [51, 52, 53] to models of gene expression, population dynamics, and competition.
@@
Here, we combine the WKB [54, 22, 25] approach of Hindes et al. [27] with the non-Markovian framework of Aquino et al. [50, 52] to study non-Markovian epidemic dynamics beyond mean field in a well-mixed setting. We compute the mean outbreak size and its full distribution within the SIR model, using gamma-distributed WTs as a prototypical example. We then analyze the SIS model, deriving the metastable mean, its full distribution, and MTE—the first passage time to the absorbing disease-free state [55, 23]. In both cases, we calculate memory kernels from general WT distributions and derive effective master equations using the final value theorem [50, 52]. Then, we apply the WKB approximation [54, 25] to obtain those key properties.
@@
In this work we focus on an example of having Markovian infection and non-Markovian recovery, with a gamma-distributed WT, while complementary cases and other WT choices are discussed in the Supplementary Information (SI) file. Here, the WT distributions satisfy
$\begin{gathered}\hskip-5.69054pt\psi_{1}(t)=\lambda_{1}e^{-\lambda_{1}t},\quad\psi_{2}(t)=(\alpha\lambda_{2})^{\alpha}t^{\alpha-1}e^{-\alpha\lambda_{2}t}/\Gamma(\alpha),\end{gathered}$ (2)
where $\Gamma\left(a,z\right)=\int_{z}^{\infty}t^{z-1}e^{-t}dt\,$ is the incomplete gamma function. Notably, the WTs means are set to equal the Markovian counterparts, $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$ respectively, where $R_{0}=\beta/\gamma$ is the basic reproductive number. The shape parameter $\alpha$ quantifies memory: small $\alpha\to 0$ yields a broad WT distribution, at $\alpha=1$ it converges to the exponential distribution and as $\alpha\to\infty$ it approaches a deterministic (delta-like) process.
To determine the outbreak size distribution within the SIR model, we write the corresponding non-Markovian master equation for the probability $P_{S,I}$ to find $S$ susceptibles and $I$ infected at time $t$, which reads [51, 52]
$\frac{\partial P_{S,I}}{\partial t}=\int_{0}^{t}\left[M_{1}(S+1,I-1,t-t^{\prime})P_{S+1,I-1}(t^{\prime})\right$.
$\left.-M_{1}(S,I,t-t^{\prime})P_{S,I}(t^{\prime})+M_{2}(S,I+1,t-t^{\prime})P_{S,I+1}(t^{\prime})\right$.
$\left.-M_{2}(S,I,t-t^{\prime})P_{S,I}(t^{\prime})\right]\,dt^{\prime}.$ (3)
In both infection and recovery, the dynamics depend on the full time history via the memory kernels $M_{i}$. The latter can be found by Laplace-transforming the master equation and using the specific structure of the WT distributions, $\psi_{i}(t)$, see SI, Sec. A. Even though the SIR dynamics has no metastability, and includes an infectious wave which quickly dies out, to explore the final outbreak distribution it suffices to look at the asymptotic late-time dependence of $M_{i}$ in the limit of $t\to\infty$ 111 This is justified by the fact that for $N\gg 1$ the epidemic duration is still much longer than the typical relaxation time of the system [59].. Below we will also provide numerical justification for this claim. Thus, we use the final value theorem (FVT): $\lim_{t\to\infty}f(t)=\lim_{u\to 0}u\tilde{f}(u)$, for the memory kernel functions, where $u$ is the Laplace variable. Defining the normalized asymptotic memory kernels: $\mathcal{M}_{i}:=(1/N)\lim_{t\to\infty}M_{i}(t)=(1/N)\lim_{u\to 0}\tilde{M}_{i}(u)$ (see details in the SI, Sec. A) and using (Large deviations in non-Markovian stochastic epidemics), we obtain the effective late-time master equation for the SIR model
@@
$+$ $\left.\mathcal{M}_{2}(S,I+1)P_{S,I+1}-\mathcal{M}_{2}(I)P_{S,I}\right],$
$\hskip-39.83385pt\text{with}\quad\mathcal{M}_{1}=R_{0}x_{i}x_{s},\quad\mathcal{M}_{2}=\frac{R_{0}x_{i}x_{s}}{\left(1+R_{0}x_{s}/\alpha\right)^{\alpha}-1}.$
This is one of the main results of this work. The asymptotic memory kernels, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are the effective infection and recovery rates under Markovian infection and non-Markovian (gamma-distributed) recovery. While for $\alpha=1$ we recover the usual rate of $x_{i}$, decreasing $\alpha$ increases the recovery rate and greatly alters the dynamics.
Having found the effective rates, we now use the WKB approximation in the spirit of [27] to compute the final outbreak size distribution; the WKB approximation is valid as long as $N\gg 1$ and the mean outbreak size is not too small. Substituting $P_{S,I}\sim e^{-N\mathcal{S}(x_{s},x_{i},t)}$ into Eq. (5) yields, in the leading order in $N\gg 1$, to a Hamilton-Jacobi equation $\partial\mathcal{S}(x_{s},x_{i},t)/\partial t+\mathcal{H}(x_{s},x_{i})\!=\!0$, where $\mathcal{S}(x_{s},x_{i},t)$ is the action function, with the Hamiltonian
@@
Here, we have defined the momenta $p_{i}=\partial\mathcal{S}/\partial x_{i}$, $p_{s}=\partial\mathcal{S}/\partial x_{s}$ in analogy to classical mechanics. In the SI, Sec. B we show that $p_{i}$ is a constant of motion, and use Hamilton’s equations to derive a relation between $p_{i}$ and final susceptible fraction $x_{s}^{*}\!=\!x_{s}(t\!\to\!\infty)$. Defining $m=e^{p_{i}}$, we find an implicit relation between $m$ and $x_{s}^{*}$
@@
This allows to compute the action function $\mathcal{S}(x_{s},x_{i},t)=\int_{0}^{t}(p_{s}\dot{x}_{s}+p_{i}\dot{x}_{i}-\mathcal{H})dt^{\prime}$ [27] (see SI, Sec. B), which reads
@@
and is a function of $x_{s}^{*}$ only via Eq. (7). Finally, the final outbreak distribution is $P(x_{s}^{*})\!\sim\!e^{-N\mathcal{S}(x_{s}^{*})}$.
To validate our theory, we run simulations using the next reaction method with simplified WT sampling [57], which in this case is accurate and faster compared to more elaborate Laplace- or non-Markovian-Gillespie schemes [41, 47]. In Fig. 1 we show examples of the final outbreak size distribution for four different $\alpha$’s. Here, theory and simulations agree well as long as the final outbreak fraction $x_{r}^{*}=1-x_{s}^{*}$ is not too close to zero.
@@
To quantify outbreak variability, we compute the standard deviation by expanding the action $\mathcal{S}$ to second order around the mean final susceptible fraction $\bar{x}_{s}^{*}$, obtained at $p_{s}=p_{i}=0$. Plugging this in Eq. (7) we get $\int_{\bar{x}_{s}^{*}}^{1}(\mathcal{M}_{2}/\mathcal{M}_{1})dx_{s}=\int_{0}^{1-\bar{x}_{s}^{*}}dx_{r}$, an implicit equation for $\bar{x}_{s}^{*}$. Using $\mathcal{M}_{i}$ from (5) yields: $\bar{x}_{s}^{*}=1-R_{0}^{-1}f(\bar{x}_{s}^{*})$, with
$f(\bar{x}_{s}^{*})\!=\!B\left[\left(1\!+\!\bar{x}_{s}^{*}R_{0}/\alpha\right)^{\alpha}\!;\!1/\alpha,0\right]\!-\!B\left[\left(1\!+\!R_{0}/\alpha\right)^{\alpha}\!;\!1/\alpha,0\right],$ (9)
@@
where $B\left[x;a,b\right]$ is the incomplete beta function. Note that for $\alpha=1$, the WT distribution becomes exponential, and we recover the known result of $1-\bar{x}_{s}^{*}\!=\!-\!\ln{\bar{x}_{s}^{*}}/R_{0}$ [8].
Using the expression for the action $\mathcal{S}$, Eq. (8), we can derive the standard deviation $\sigma$ by performing a Gaussian approximation on $P(x_{s}^{*})$ around $\bar{x}_{s}^{*}$. This yields $\sigma\simeq\left|-N\mathcal{S}^{\prime\prime}(\bar{x}_{s}^{*})\right|^{-1/2}$, which can be directly compared to simulation results. Figure 2 shows the effect of varying the shape parameter $\alpha$ for gamma-distributed recovery with exponential infection: as $\alpha$ decreases, the mean outbreak size drops (since typical recovery times are shorter than in the Markovian case) while the standard deviation increases. We also see that the WKB approximation breaks down when the mean outbreak size nears zero.
@@
Figure 1: Outbreak size distributions at $\alpha=0.65$, $0.8$, 11 and $10$ in (a)–(d), for Markovian infection and gamma-distributed recovery, $N=5000$, $R_{0}=1.5$, and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (solid line). In all panels $I(0)=1$, and outbreak fractions $10^{-2}$ were omitted.
@@
Figure 2: Normalized mean $x_{r}^{*}$ (a) and normalized standard deviation $\sigma$ (b) versus $\alpha$. Insets show those quantities versus $\tilde{\alpha}\equiv(\alpha-\alpha_{bif})/\alpha_{bif}$, where at $\alpha_{bif}$ the mean vanishes. Here infection is exponential and recovery is gamma-distributed, $N=5000$, $R_{0}=1.5$, $\alpha_{bif}\simeq 0.5$, and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (solid line). In all panels $I(0)=1$, and outbreak fractions $10^{-2}$ were omitted.
@@
To quantify the effect of non-Markovian recovery on the endemic state $x_{i}^{*}$, we multiply Eq. (10) by $I$ and sum over all $I$. This yields the modified rate equation $\langle\dot{x}_{i}\rangle=\mathcal{M}_{1}(\langle x_{i}\rangle)-\mathcal{M}_{2}(\langle x_{i}\rangle)$, with $x_{i}^{*}$ found by solving $\mathcal{M}_{1}(x_{i}^{*})=\mathcal{M}_{2}(x_{i}^{*})$. Using (5) with $x_{s}=1-x_{i}$ gives
$x_{i}^{*}=1-(\alpha/R_{0})(2^{1/\alpha}-1).$ (11)
At $\alpha=1$ (exponential WTs), we recover the Markovian result $x_{i}^{*}=1-1/R_{0}$ with a bifurcation at $R_{0}=1$. For $\alpha\neq 1$, the bifurcation shifts to $R_{bif}=\alpha(2^{1/\alpha}-1)$: it moves above 1 for $\alpha<1$ and below 1 for $\alpha>1$. Consequently, we find that an endemic state can persist even when each infected transmits to fewer than one other individual, or conversely, a highly infectious disease may still die out exponentially fast. Note that the relaxation time of the dynamics $\tau_{r}=\left[\mathcal{M}_{2}{}^{\prime}\left(x_{i}^{*}\right)-\mathcal{M}_{1}{}^{\prime}\left(x_{i}^{*}\right)\right]^{-1}$ can also be computed. Using Eq. (5) we obtain $\tau_{r}=2^{1/\alpha-1}/\left\{\alpha\left(2^{1/\alpha}\!-\!1\right)\left[R_{0}\!-\!\alpha\left(2^{1/\alpha}\!-\!1\right)\right]\right\}$, which coincides with the known result of $\tau_{r}=1/(R_{0}-1)$ at $\alpha=1$.
We now analyze the effective master equation (10) to determine the QSD of the metastable state and the MTE under non-Markovian recovery. Assuming the metastable state decays slowly, we set $P_{I}(t)=\pi(x_{i})e^{-t/\tau_{ext}}$, where $\pi(x_{i})$ is the QSD and $\tau_{ext}$ is the exponentially large MTE. We now apply the WKB approximation, $\pi(x_{i})\sim e^{-N\mathcal{S}(x_{i})}$, where $\mathcal{S}(x_{i})$ is the action function. Expanding to leading order in $N\gg 1$, and using Eq. (5) we obtain $\mathcal{S}^{\prime}(x_{i})=-\ln[\mathcal{M}_{1}(x_{i})/\mathcal{M}_{2}(x_{i})]=-\ln[(1+R_{0}(1-x_{i})/\alpha)^{\alpha}-1]$, which provides the QSD $\pi(x_{i})$

### v2

@@ Abstract
We develop a framework for non-Markovian, well-mixed SIR and SIS models beyond mean field, utilizing the continuous-time random walk formalism. Using a gamma distribution for the infection and recovery inter-event times as a test case, we derive asymptotical late-time master equations with effective memory kernels and obtain analytical predictions for the final outbreak size distribution in the SIR model, and quasistationary distribution and disease lifetime in the SIS model. We show that varying the width of the inter-event time distribution can greatly alter the outbreak size distribution or the disease lifetime. We also show that rescaled Markovian models may fail to capture fluctuations in the non-Markovian case. Overall, our analysis, confirmed against numerical simulations, paves the way for studying large deviations in structured populations on degree-heterogeneous networks.
@@
Introduction. The challenge of modeling and predicting pandemics has motivated extensive development of deterministic and stochastic epidemiological models [4, 35, 2, 45, 19, 26, 33, 46]. Their importance was underscored during COVID-19 [68, 52], when forecasts guided policies worldwide. These models typically divide populations into disease-status compartments. The SIS model [25] with susceptible and infected compartments, describes infections with short-lived immunity (e.g., influenza, or the common cold), while the SIR model [46] adds a recovered compartment to capture long-term immunity, which is relevant for diseases such as measles, smallpox, polio and COVID-19. Generalized models include reinfection (SIRS), an exposed compartment (SEIR) [26], time-varying rates [6, 55], or structured populations, where network-based approaches [50, 32, 41, 44, 34, 31, 15] capture heterogeneity in population structure. Mean-field quantities have been computed for both models: the mean outbreak size and epidemic threshold in the SIR model, and endemic state and relaxation dynamics in the SIS model [25, 19, 2, 50]. Stochastic dynamics have also been explored, including the outbreak-size distribution for the SIR model and quasistationary distribution (QSD) and disease mean time to extinction (MTE) for the SIS model [48, 21, 55, 47, 7, 47, 30, 8, 28, 27, 29, 39, 38].
A key limitation of most approaches is the Markovian assumption [33, 65, 56], which states that the system dynamics depend solely on the present state, and are independent of prior history. Under this assumption, both infection and recovery periods are exponentially distributed. However, empirical evidence suggests that transmission and recovery rates are often explicitly time-dependent, and consequently, the corresponding waiting-time (WT) distributions are more accurately described by non-exponential distributions, such as log-normal, gamma, or Weibull [10, 23, 66]. As a result, infection and recovery dynamics are generally non-Markovian, rendering analytical frameworks based on Markovian assumptions often inadequate for describing such processes.
@@
Most studies investigating non-Markovian epidemic processes on various network topologies, have focused on the mean-field aspects of the dynamics [17, 60, 13, 36, 53, 37, 56, 54, 43, 57]. These studies have shown that, even when average rates are unchanged, non-exponential infection and recovery times can drastically affect disease spread, prevalence, and overall epidemic bahavior. Conversely, certain features of non-Markovian dynamics can still be approximated by rescaling Markovian rates, see, e.g., Refs. [57, 22]. In this context, Boguñá et al. [13] proposed a generalized Gillespie algorithm applicable to non-exponential infection and recovery processes, which was later refined into the more efficient Laplace Gillespie method by Masuda et al. [43]. Notably, while the role of noise in non-Markovian epidemics has also been studied in both SIR [11, 58, 18, 67] and SIS [17] models, a systematic analysis of the interplay between non-Markovianity and demographic noise on large deviations in these models, remains lacking.
@@
Recent progress has extended non-Markovian frameworks to chemical and biological systems, using the continuous-time random walk formalism with non-exponential WTs [5]. This was later adapted in [62, 61, 63] to models of gene expression, population dynamics, and competition. Here, we combine this framework with the WKB approach [20, 47, 8] of Hindes et al. [27] to study non-Markovian epidemic dynamics beyond mean field in a well-mixed setting. We calculate the memory kernels from general WTs and derive the effective late-time master equation [5, 61]. We then compute the mean outbreak size and its full distribution within the SIR model, using gamma-distributed WTs as a prototypical example. We then analyze the SIS model, deriving the metastable mean, its full distribution (QSD), and MTE—the first passage time to the absorbing, disease-free state [51, 7]. Finally, we test our formalism using empirical WTs for both infection and recovery.
It is important to note that, in this work we focus on well-mixed populations, i.e., fully-connected networks, in which all individuals are considered identical. As such, for simplicity, we consider only two global reactions: infection and recovery, both with generic WT distributions, whose mean depends on the system’s state. While this constitutes an oversimplified representation of realistic scenarios, it turns out that our model can reproduce, at least qualitatively, key features of large deviations in non-Markovian epidemic dynamics on complex networks.
@@
Initially we focus on the case of non-Markovian, gamma-distributed infection and Markovian recovery [60, 13], while other WT choices are discussed in the Supplementary Information (SI). The WT distributions satisfy
$\begin{gathered}\hskip-8.53581pt\psi_{1}(t)=(\alpha\lambda_{1})^{\alpha}t^{\alpha-1}e^{-\alpha\lambda_{1}t}/\Gamma(\alpha),\quad\psi_{2}(t)=\lambda_{2}e^{-\lambda_{2}t},\end{gathered}$ (2)
where $\Gamma\left(z\right)=\int_{0}^{\infty}t^{z-1}e^{-t}dt\,$ is the gamma function. Notably, the definition of the gamma distribution here is such that the shape parameter $\alpha$ does not affect the mean. The WTs means are set to equal the Markovian counterparts, $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$ respectively, where $R_{0}=\beta/\gamma$ is the basic reproductive number. The shape parameter $\alpha$ controls the tail of the WT distribution, and as such, can be effectively viewed as a form of memory, where lower values of $\alpha$ entail stronger memory (broader distribution) while higher values of $\alpha$ entail weaker memory (narrower distribution).
To determine the outbreak size distribution within the SIR model, we write the corresponding non-Markovian master equation for the probability $P_{S,I}$ to find $S$ susceptibles and II infected at time $t$, which reads [62, 61]
$\hskip-8.53581pt\frac{\partial P_{S,I}}{\partial t}=\int_{0}^{t}\left[M_{1}(S+1,I-1,t-t^{\prime})P_{S+1,I-1}(t^{\prime})\right$.
$\hskip-8.53581pt\left.-M_{1}(S,I,t-t^{\prime})P_{S,I}(t^{\prime})+M_{2}(S,I+1,t-t^{\prime})P_{S,I+1}(t^{\prime})\right$.
$\hskip-8.53581pt\left.-M_{2}(S,I,t-t^{\prime})P_{S,I}(t^{\prime})\right]\,dt^{\prime}.$ (3)
In both infection and recovery, the dynamics depend on the full time history via the memory kernels $M_{i}$. The latter can be found by Laplace-transforming the master equation and using the specific structure of the WT distributions, $\psi_{i}(t)$, see SI, Sec. A. Even though the SIR dynamics has no metastability, and includes an infectious wave which quickly dies out, to explore the final outbreak distribution it suffices to look at the asymptotic late-time dependence of $M_{i}$ in the limit of $t\to\infty$ [59]  [^1] This is justified since for $N\gg 1$ the epidemic duration is still much longer than the system’s relaxation time [57].. Below we will also provide numerical justification for this claim. Thus, we use the final value theorem: $\lim_{t\to\infty}f(t)=\lim_{u\to 0}u\tilde{f}(u)$, for the memory kernel functions, where $u$ is the Laplace variable. Defining the normalized asymptotic memory kernels: $\mathcal{M}_{i}:=(1/N)\lim_{t\to\infty}M_{i}(t)=(1/N)\lim_{u\to 0}\tilde{M}_{i}(u)$ (see details in the SI, Sec. A) and using (Large deviations in non-Markovian stochastic epidemics), we obtain the effective late-time master equation for the SIR model
@@
$+$ $\left.\mathcal{M}_{2}(S,I+1)P_{S,I+1}-\mathcal{M}_{2}(S,I)P_{S,I}\right],$
$\hskip-39.83385pt\text{with}\quad\mathcal{M}_{1}=\frac{x_{i}}{\left[1+1/(R_{0}x_{s}\alpha)\right]^{\alpha}-1},\quad\mathcal{M}_{2}=x_{i},$
where $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, are the effective infection and recovery rates under non-Markovian (gamma-distributed) infection and Markovian recovery. For $\alpha=1$ (exponential infection), we recover the Markovian rate $\mathcal{M}_{1}=R_{0}x_{s}x_{i}$, whereas decreasing $\alpha$ increases the infection rate and greatly alters the dynamics. Notably, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ can also be found for other choices of WTs, see SI, Sec. B.
Having found the effective rates, we now use the WKB approach as in [27] to compute the final outbreak-size distribution. Substituting $P_{S,I}\sim e^{-N\mathcal{S}(x_{s},x_{i},t)}$ into Eq. (5) yields, in the leading order in $N\gg 1$  [^2] Note that, the leading-order WKB approximation here is valid as long as the total action, $N\mathcal{S}$, is large [20, 7, 8]., a Hamilton-Jacobi equation $\partial\mathcal{S}(x_{s},x_{i},t)/\partial t+\mathcal{H}(x_{s},x_{i})\!=\!0$, where $\mathcal{S}(x_{s},x_{i},t)$ is the action function, with the Hamiltonian
@@
Here, we have defined the momenta $p_{i}=\partial\mathcal{S}/\partial x_{i}$, $p_{s}=\partial\mathcal{S}/\partial x_{s}$ in analogy to classical mechanics. In the SI, Sec. C we show that $p_{i}$ is a constant of motion, and use Hamilton’s equations to derive a relation between $p_{i}$ and final susceptible fraction $x_{s}^{*}\!=\!x_{s}(t\!\to\!\infty)$. Defining $m=e^{p_{i}}$, we find an implicit relation between $m$ and $x_{s}^{*}$
@@
This allows to compute the action function $\mathcal{S}(x_{s},x_{i},t)=\int_{0}^{t}(p_{s}\dot{x}_{s}+p_{i}\dot{x}_{i}-\mathcal{H})dt^{\prime}$ [27] (see SI, Sec. C), which reads
@@
and is a function of $x_{s}^{*}$ only via Eq. (7). Finally, the final outbreak distribution is given by $P(x_{s}^{*})\!\sim\!e^{-N\mathcal{S}(x_{s}^{*})}$.
@@
To quantify outbreak variability, we compute the standard deviation by expanding the action $\mathcal{S}$ to second order around the mean final susceptible fraction $\bar{x}_{s}^{*}$, obtained at $p_{s}=p_{i}=0$. Plugging this in Eq. (7) we get $\int_{\bar{x}_{s}^{*}}^{1}(\mathcal{M}_{2}/\mathcal{M}_{1})dx_{s}\!=\!\int_{0}^{1-\bar{x}_{s}^{*}}\!\!dx_{r}$, an implicit equation for $\bar{x}_{s}^{*}$. Using $\mathcal{M}_{i}$ from (5) yields: $1-\bar{x}_{s}^{*}=(2\alpha R_{0})^{-1}f(\bar{x}_{s}^{*})$, with
$\color[rgb]{0,0,0}f(\bar{x}_{s}^{*})\!=\!\mathrm{B}\!\left(\!\frac{\alpha R_{0}}{1\!+\!\alpha R_{0}};\!1\!-\!\alpha,-1\!\right)\!-\!\mathrm{B}\!\left(\!\frac{\alpha R_{0}\bar{x}_{s}^{*}}{1\!+\!\alpha R_{0}\bar{x}_{s}^{*}}\!;1\!-\!\alpha,-1\!\right)\!,$} (9)
@@
where $\mathrm{B}(x;a,b)\!=\!\!\int_{0}^{x}\!t^{1\!-\!a}(1\!-\!t)^{1\!-\!b}dt$ is the incomplete beta function. Here, for $\alpha\!=\!1$, we recover the known mean-field result of $1-\bar{x}_{s}^{*}\!=\!-\!\ln{\bar{x}_{s}^{*}}/R_{0}$ [46].
Using Eq. (8) for the action $\mathcal{S}$, we can derive the standard deviation $\sigma$ by performing a Gaussian approximation on $P(x_{s}^{*})$ around $\bar{x}_{s}^{*}$, yielding $\sigma\!\simeq\!\left|N\mathcal{S}^{\prime\prime}(\bar{x}_{s}^{*})\right|^{-1/2}$ [27]. Defining $\xi(\bar{x}_{s}^{*})\!=\!\mathcal{M}_{2}(\bar{x}_{s}^{*})/\mathcal{M}_{1}(\bar{x}_{s}^{*})$ and $I(\bar{x}_{s}^{*})=\int_{\bar{x}_{s}^{*}}^{1}\xi(z)^{2}dz$, we find (see SI, Sec. C)
$\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}=\left[\xi(\bar{x}_{s}^{*})-1\right]^{2}/\left[1-\bar{x}_{s}^{*}+I(\bar{x}_{s}^{*})\right],$ (10)
where using Eq. (5), $\xi(\bar{x}_{s}^{*})\!=\![1\!+\!1/(R_{0}\alpha\bar{x}_{s}^{*})]^{\alpha}\!-\!1$, and $I(\bar{x}_{s}^{*})$ can be expressed by beta functions.
To validate our theory, we ran simulations using the next reaction method with simplified WT sampling  [^3] In each realization, we sample the WTs for both infection and recovery processes from the given WT distributions, execute the reaction that occurs first, advance the system time accordingly, update the distributions’ mean rates, reset the timers (as done in Ref. [43] when the WT means are state-dependent), and repeat until either disease extinction occurs or the maximum time is reached., see [3], which in this case is accurate and faster compared to the algorithms in Refs. [13, 43]. Figures 1 and 2 show results for the non-Markovian SIR model with gamma-distributed infection and exponential recovery. In Fig. 1 we show examples of the final outbreak size distribution for four different $\alpha$’s  [^4] In these simulations, we performed a sufficiently large number of realizations so that, even after excluding those with a final outbreak fraction $10^{-2}$, we still retained the desired number of samples, see Figs. 1 and 2. This allowed us to probe the distribution around the mean without the influence of near-immediate extinction events.. Here, theory and simulations agree well as long as the action is large [20, 7, 8, 27]. In Fig. 2 we show the mean outbreak fraction and its variance as function of $\alpha$. As $\alpha$ increases, typical infection times becomes longer, and the mean drops, while the standard deviation increases. We also see that the WKB approximation loses its accuracy as the mean outbreak size approaches zero.
@@
Figure 1: Outbreak size distributions for the SIR model, with $\alpha=0.5$, 11, $1.5$ and 22 in (a)–(d), for gamma-distributed infection and Markovian recovery, $N=5000$, $R_{0}=1.5$, and $10^{5}$ runs per $\alpha$. Simulations (symbols) are compared with theory (solid line). In all panels $I(0)=1$, and simulations resulting in outbreak fractions $10^{-2}$ were omitted, see [61].
@@
Figure 2: Normalized mean $x_{r}^{*}$ (a) and normalized standard deviation $\sigma$ (b) versus $\alpha$. Here infection is gamma-distributed and recovery is exponential, $N=5000$, $R_{0}=1.5$ and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (solid line). In all panels $I(0)=1$, and simulations resulting in outbreak fractions $10^{-2}$ were omitted, see [61].
@@
To quantify the effect of non-Markovian infection on the endemic state $x_{i}^{*}$, we multiply Eq. (11) by II and sum over all II. This yields the modified rate equation $\langle\dot{x}_{i}\rangle=\mathcal{M}_{1}(\langle x_{i}\rangle)-\mathcal{M}_{2}(\langle x_{i}\rangle)$, with $x_{i}^{*}$ found by solving $\mathcal{M}_{1}(x_{i}^{*})=\mathcal{M}_{2}(x_{i}^{*})$. Using (5) with $x_{s}=1-x_{i}$ gives
$\color[rgb]{0,0,0}x_{i}^{*}=1-[\alpha R_{0}(2^{1/\alpha}-1)]^{-1}.$} (12)
At $\alpha=1$ (exponential WTs), we recover the Markovian result $x_{i}^{*}=1-1/R_{0}$. However, at $\alpha\neq 1$, one can define a new effective $R_{0}$, $R_{0}^{\mathrm{eff}}=\alpha R_{0}(2^{1/\alpha}-1)$, such that $x_{i}^{*}=1-1/R_{0}^{\mathrm{eff}}$. This means that, in order to reproduce the effect of non-Markovian infection, one can instead take Markovian infection and change the infection rate $\beta\to\beta\alpha(2^{1/\alpha}-1)$ such that at $\alpha\to 0$, the Markovian infection rate should tend to $\infty$, while at $\alpha\to\infty$, the infection rate should tend to $\beta\ln 2$. Consequently, we find that, at $\alpha<1$ an endemic state can persist even when each infected transmits to fewer than one other individual, or in other words, a subcritical epidemic with $R_{0}<1$ may still reach a nonzero endemic state. Note that the relaxation time of the dynamics $\tau_{r}=\left[\mathcal{M}_{2}{}^{\prime}\left(x_{i}^{*}\right)-\mathcal{M}_{1}{}^{\prime}\left(x_{i}^{*}\right)\right]^{-1}$ can also be computed. Using Eq. (5) we obtain $\tau_{r}=2^{1/\alpha-1}/\left\{\alpha\left(2^{1/\alpha}\!-\!1\right)\left[R_{0}\alpha\left(2^{1/\alpha}\!-\!1\right)\!-\!1\right]\right\}$, which coincides with the known result of $\tau_{r}=1/(R_{0}-1)$ at $\alpha=1$.
We now analyze the effective master equation (11) to determine the QSD of the metastable state and the MTE under non-Markovian recovery. Assuming the metastable state decays slowly, we set $P_{I}(t)=\pi(x_{i})e^{-t/\tau_{ext}}$, where $\pi(x_{i})$ is the QSD and $\tau_{ext}$ is the exponentially large MTE. We now apply the WKB approximation, $\pi(x_{i})\sim e^{-N\mathcal{S}(x_{i})}$, where $\mathcal{S}(x_{i})$ is the action function. Expanding to leading order in $N\!\gg\!1$, and using (5) we find $\mathcal{S}^{\prime}(x_{i})=\ln[\mathcal{M}_{2}(x_{i})/\mathcal{M}_{1}(x_{i})]=\ln\{[1+1/(R_{0}\alpha(1\!-\!x_{i}))]^{\alpha}-1\}$, which provides the QSD, $\pi(x_{i})$